OpenStudy (anonymous):

Let f(x)=2x^3-5x^2 and g(x)=2x-1 find ( f * g)(-2)

OpenStudy (anonymous):

(f*g)(x)=f(x)*g(x)=(2x^3-5x^2)(2x-1) am i right?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

put x=-2 what will u get?